OpenStudy (anonymous):

Which of the following sets would have a graph that is a closed circle on -4 and an arrow to the left? {x | x ∈ R, x < -4} {x | x ∈ R, x ≤ -4} {x | x ∈ R, x > -4} {x | x ∈ R, x ≥ -4}

OpenStudy (anonymous):

closed circle mean that the value will be included in the set so you will need either\[\le or \ge\]

OpenStudy (anonymous):

{x | x ∈ R, x ≤ -4}

OpenStudy (anonymous):

To the left means you want values that are less than that value

OpenStudy (anonymous):

so its the bottom one

OpenStudy (anonymous):

no marvin is correct, remember your symbols \[\le\] is less than \[\ge\] is greater than

OpenStudy (anonymous):

I mean less than or equal and greater than or equal...

OpenStudy (anonymous):

thanks and ya i forgot ;/

OpenStudy (anonymous):

:)